The magnetic bearing of line PQ is S40E and the magnetic declination is 8, 5'E what is the true bearing of the line?
Answer #1

The W.C.B of line PQ =180 - 40 =40 degrees

True bearing of PQ = Magnetic bearing +magnetic declination E

                           = 140+8,5'

True bearing of PQ =148 degrees 5'